# The Reliability of Remotely Piloted Aircraft System Performance under Communication Loss and Latency Uncertainties

### v1

@@ 6 Conclusions
This study examines how latency and communication availability influence the flight performance of a Remotely Piloted Aircraft System (RPAS), modeled on F-16 dynamics, during a static waypoint-following task. We begin by reviewing relevant aviation communication standards, including ICAO’s Required Communication Performance (RCP) and JARUS’s Required Link Performance (RLP), and then derive their corresponding mathematical formulations. In addition, we propose a new metric, communicability, that collectively represents availability, continuity, and transaction time.
@@ 6.2 Future Studies
    The current work focuses on static waypoint-following tasks. However, a significant portion of interests lies in flight performance within highly dynamic environments, typically involving multiple moving agents (e.g., collision avoidance, separation violations, tracking). Performance in such dynamic scenarios can be substantially more complex than those examined here.

### v2

@@ 6 Conclusions
This study examines how latency and communication availability influence the flight performance of a Remotely Piloted Aircraft System (RPAS), modeled on flight dynamics, during a static waypoint-following task. We begin by reviewing relevant aviation communication standards, including ICAO’s Required Communication Performance (RCP) and JARUS’s Required Link Performance (RLP), and then derive their corresponding mathematical formulations. In addition, we propose a new metric, communicability, that collectively represents availability, continuity, and transaction time.
@@ 6.2 Future Studies
    The current work focuses on static waypoint-following tasks. However, a significant portion of interests lies in flight performance within highly dynamic environments, typically involving multiple moving agents (e.g., collision avoidance, separation violations, tracking). Performance in such dynamic scenarios can be substantially more complex than those examined here. In system-of-systems contexts, system importance measures could help identify the most vulnerable subsystems and guide design improvements for resilience (Uday et al., 2019).